Case 1 – New data set: ABP = 90/60 mm Hg, HR 145 Beats/min, cardiac output = 3 l/min, CVP = 0 mm Hg, PAP = 30/10 mm Hg, PCWP = 2 mm Hg. Calculate the SV and fill in the table below and answer the question below.

Mean arterial pressure (MAP)

mmHg

Systemic vascular resistance (SVR)

dynes

Mean pulmonary artery pressure (PA mean)

mmHg

Pulmonary vascular resistance (PVR)

dynes

Stroke volume (SV)

ml

Regarding Case 1. The best treatment for this patient is probably: (choose one and explain your reason).

diuretics

aortic balloon pump

intravenous fluid

administration vasoconstrictors
Answer #1

Answer :- Given that,

ABP = 90/60 mm Hg, HR 145 Beats/min, cardiac output = 3 l/min, CVP = 0 mm Hg, PAP = 30/10 mm Hg, PCWP = 2 mm Hg,

Mean arterial pressure (MAP)

70                mm Hg

Systemic vascular resistance (SVR)

1867                dynes

Mean pulmonary artery pressure (PA mean)

70                   mm Hg

Pulmonary vascular resistance (PVR)

1813                        dynes

Stroke volume (SV)

21                      ml

The best treatment for patient is Intravenous fluid administration. It will help the patient to normalize the blood pressure, vasocontrictors cannot be used it will prompt increment in precise blood stream yet my happen on particular tissue type cauaica a restricted decrease in blood stream

Diuretics will bring about additional bringing down in Blood weight

The heart yield is lower than ordinary principally in view of after load is expanded.